Suppose that on January 2, 2018, the U.S. Treasury auctioned the following four T-Bills: 28-Day 91-Day...

  1. Suppose that on January 2, 2018, the U.S. Treasury auctioned the following four T-Bills:

28-Day

91-Day

182-Day

270-Day

Issue Date

1/2/2018

1/2/2018

1/2/2018

1/2/2018

Maturity Date

1/30/2018

4/3/2018

7/3/2018

9/29/2018

Face Value per $100

100

100

100

100

Price per $100

99.8

99.2

98.3

97.5

  1. Determine the bank discount rate for each T-Bill using formula 10-3 given in the chapter and the DISC function.
  1. Determine the bond equivalent yield for each T-Bill using the formulas given in the chapter and the YIELDDISC function.
Answer #1

As it is already mentioned that we have to use DISC and YIELDDISC functions. So here I am explaining both the functions before proceeding

DISC Function - this function returns the discount rate of security using inputs such as settlement date, maturity date, security price per $100 face value, security redemption value per $100 face value, and basis which is optional to use day count convention

=DISC (settlement, maturity, pr, redemption, [basis])

YIELDDISC - this function gives an annual yield for a discounted security using the same inputs like DISC function. The difference is that in YIELDDISC the return is annualized

Inputs for 28-day bond

Settlement date = T+1 = 1/3/2018

Maturity date = 1/ 30/ 2018

Redemption = 100

Price = 99.8

Bank discount rate = 2.66%

Bank equivalent yield = 2.67%

Inputs for 91-day bond

Settlement date = T+1 = 1/3/2018

Maturity date = 4/ 3/ 2018

Redemption = 100

Price = 99.2

Bank discount rate = 3.2%

Bank equivalent yield = 3.23%

Inputs for 182-day bond

Settlement date = T+1 = 1/3/2018

Maturity date = 7/ 3/ 2018

Redemption = 100

Price = 98.3

Bank discount rate = 3.4%

Bank equivalent yield = 3.46%

Inputs for 270-day bond

Settlement date = T+1 = 1/3/2018

Maturity date = 29/9/ 2018

Redemption = 100

Price = 97.5

Bank discount rate = 3.38%

Bank equivalent yield = 3.47%